OpenStudy (anonymous):

Solve the system algebraically. 2x + 5y = 10 2x + 3y = 6

OpenStudy (anonymous):

do u try it yet ??

Parth (parthkohli):

Use elimination. Subtract both the equations, and then find the value of x.

OpenStudy (callisto):

2x + 5y = 10 -(1) 2x + 3y = 6 -(2) (1)-(2) 5y-3y = 10-6 2y = 4 What do you get for y?

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 2y = 4 }\) \(\Large \color{purple}{\rightarrow y = ? }\) \(\Large \color{purple}{\rightarrow 2x + 5(?) = 10 }\)

Parth (parthkohli):

Now you can find the value of both.

OpenStudy (lgbasallote):

lol q marks

OpenStudy (anonymous):